[property tax revenue has dropped the past four years. If tax revenue in 2011 was $50,000,000 less than twice the revenue in 2014 and revenue dropped $24,000,000 from 2011 to 2014, find the revenue for each year?]
Answer by TimothyLamb(4379)

x = revenue 2011
y = revenue 2014
---
x = 2y - 50000000
x - y = 24000000
---
put the system of linear equations into standard form
---
x - 2y = -50000000
x - y = 24000000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 98000000
y = 74000000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php